Previous | Calendar |
September |
2009 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
30JD242 |
31JD243 |
1JD244Site Visit (SpB, DA, DB, AT, T): Details , Photos (27) Span cylinder at 1200 psi. N2 at 300 psi. removed loaner LI-6262 (sn IRG3-224); re-installed our re-calibrated LI-6262 (sn IRG3-0638) meet Joe Lanzoni at 8:40 at MRS. Smoke/Haze from fires (california?) is evident. |
2JD245 |
3JD246 |
4JD247(Full "Corn" Moon) |
5JD248 |
6JD249 |
7JD250Site Visit (SpB): Details , Photos (17) Span cylinder at 1150 psi. N2 at 100 psi --> 2000 psi from 10:00-10:30 MST, switched N2 cylinder at the tower (new cylinder=CC264924 (2000 psi); old cylinder=CC199722 (100 psi)) at around 12:00 MST, reset GFI outlet in the USGS red shed (LI-7000) from 10:00-10:35 MST, Changed profiler (li-6251) dessicant Cool and a bit breezy. Aspens are starting to turn yellow. CU holiday (Labor Day) |
8JD251 |
9JD252 |
10JD253 |
11JD254Site Visit (SpB, DA): Details , Photos (19) Span cylinder at 1050 psi. N2 at 1880 psi at 11:00 MST, re-installed krypton hygrometer replaced gelman filters for two tower inlets reconfigured the mounting for the star-dot camera (need to add guy wires?) Fairly calm and cool. Nice day to work on the tower. |
12JD255Cold and Windy (some snow, but no accumulation). |
13JD256 |
14JD257 |
15JD258 |
16JD259Site Visit (SpB, JK): Details , Photos (17) Span cylinder at 1000 psi. N2 at 1750 psi Partly overcast with a few sunbreaks & afternoon clearing. Feels like autumn has arrived (but not much wind). |
17JD260 |
18JD261 |
19JD262 |
20JD263 |
21JD264First significant snowfall of the season. |
22JD265Site Visit (SpB, CG, NT, LY): Details , Photos (13) Span cylinder at 950 psi. N2 at 1500 psi mounted boom for EOL csat-3 (sn 0674?) changed the russter2 UPS battery reconfigured EOL T/RH sensor Snowing and calm in the morning. Snowy and windy in the afternoon. Fall Equinox |
23JD266 |
24JD267 |
25JD268 |
26JD269 |
27JD270Site Visit (SpB): Details , Photos (24) Span cylinder at 820 psi. N2 at 1490 psi mounted EOL csat-3 (sn 0674) at 11:00 MST, started ops52/ops53; collecting eol csat3, accelerometer data no longer saved (after a few adjustments). started running "fast_10.csi" in fast cr23x logger moved EOL T/RH sensor up to 4.6m A bit breezy in the morning, but calm and warm most of the day (very dry). |
28JD271 |
29JD272 |
30JD273 |
1JD274 |
2JD275Site Visit (SpB): Details , Photos (25) Span cylinder at 750 psi. N2 at 1260 psi at 9:30 MST, rebooted quacker (data system back up) from 11:20-11:47 MST, Changed profiler (li-6251) dessicant removed the KZ PAR-LITE sensor (PAR2.up) from the CNR logger moved EOL T/RH sensor up to 8m re-connected noahs snow depth logger and confirmed it was running Very windy in the morning, but calm in the afternoon (very dry). |
3JD276 |